[Hallc_running] [New Logentry] Hall C RC update
brads at jlab.org
brads at jlab.org
Tue Jul 2 08:40:08 EDT 2019
Short answer is that we'll be down for at least a couple days.
To recap last night, we basically lost all power to everything inside the accelerator fence for 1.5 hours last night due the Dominion Warwick substation tripping off. I gather Dominion is working the problem but does not know the root cause yet.
Teams and systems responded well to the event and there seem to be no major knock-on failures at this point, but there is a long punch list. A relief valve on the CHL was damaged (won't reseal) and requires repair and there are some controls issues with a supporting cooling tower that is impeding progress at the moment. Some helium was lost in the cryomodules, but nothing too serious. Steve Suhring (OPS) has a nice summary of the overall status here: https://logbooks.jlab.org/entry/3697566
On our end things seem to be generally OK -- we won't hold up on the program anyway. The target has warmed up and we lost LHe in HMS Q1 and SHMS BH, Q2, and D. The ESR is online so we will coordinate refilling our magnets along with Hall A. There are a few other items to address (dead operator computer, some IOCs are still offline, HV controls has a communications glitch) but again, nothing too serious. We should try to get back to taking cosmics by the end of today.
** All shifts are cancelled today (Tues), Wed, and Thurs OWL. I'll leave Thurs Day, Swing shifts on standby for the time being. **
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3697809
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running